# PhysBrain: Human Egocentric Data as a Bridge from Vision Language Models to Physical Intelligence

### v1

@@ 3.1.3 Quality Assurance and Validation Logic
Open-ended generation easily produces errors that are harmful for training supervision. Common failures include references to non-visible hands, incorrect temporal ordering, and under-specified placeholders. We therefore introduce a deterministic rule checker as a validation gate. Samples that fail validation are rejected and sent back for regeneration with structured error messages that indicate the violated constraint.
The checker applies three types of constraints. Evidence grounding requires that all mentioned actions, hands, and contact states are supported by the clip frames. Egocentric consistency enforces the correct hand references and prohibits mentions of unseen limbs or contradictory assignments. Mode-specific temporal logic requires explicit temporal connectors for temporal-sensitive modes and verifies that the described order matches the clip timeline. The generation–validation loop repeats until all constraints are satisfied, producing supervision that is consistent, temporally coherent, and suitable for embodied learning.
3.1.4 Structured Egocentric Supervision Output
Samples that satisfy all validation constraints are retained and compiled into the egocentric VQA supervision dataset. Each entry records the sampled frames, the selected VQA mode and template, the generated question–answer pair, and the validation outcome. This design ensures traceability and reproducibility.
The dataset produced by the proposed data engine offers structured and logic-verified supervision that encodes action organization and hand–object interaction, completing the translation of egocentric video data into reliable training signals for egocentric planning and interaction reasoning.
Figure 3: Overview and Data Distribution Statistics of E2E-3M dataset.
@@ 3.2.1 Data Sources and Domain Coverage
The proposed Egocentric2Embodiment Translation Pipeline is applied to large-scale human egocentric video corpora collected across three complementary domains: household, factory, and laboratory environments, as shown in Fig.3(a). Collectively, these corpora comprise thousands of hours of egocentric video and capture substantial variation in environmental context, object composition, and interaction patterns.
Specifically, Ego4D represents open-world household activities and provides extensive geographic and contextual diversity. BuildAI captures real industrial workflows, emphasizing procedural regularity and dense hand visibility in factory environments. EgoDex focuses on laboratory settings and offers high-resolution egocentric manipulation sequences with fine-grained interaction cues. These sources differ systematically in spatial layout, object distribution, and task structure. The aggregation yields the Egocentric2Embodiment dataset with complementary coverage across the space of egocentric embodied experience.
@@ 3.2.2 Diversity Analysis
To evaluate whether the dataset provides sufficiently rich supervision for embodied planning and interaction, we analyze diversity along two interpretable axes: object coverage and action (verb) coverage in Fig.3(b). These dimensions correspond to what entities are involved in interactions and how those interactions are performed.
Object coverage measures how many distinct objects appear in the dataset annotations. It reflects the breadth of perceptual and interactional contexts captured. For each domain $s$, the object coverage is calculated as:
$\mathrm{ObjectDiv}(s)=\frac{|\mathcal{V}^{\text{noun}}_{s}|}{T^{\text{noun}}_{s}}\times 1000,$ (1)
where $\mathcal{V}^{\text{noun}}_{s}$ is the number of unique noun lemmas and $T^{\text{noun}}_{s}$ is the total noun token count in domain $s$. $\mathrm{ObjectDiv}$ values are grouped into four descriptive ranges: low (<200<200), medium (200200300300), high (300300350350), and very high ($\geq 350$).
As shown in Fig.3(b), Household data falls into the high to very high range, showing broad object diversity typical of open environments. Lab data falls in the medium range, consistent with a more limited set of experimental objects. Factory data shows low to medium coverage, reflecting repeated use of domain-specific parts and tools. These domain differences confirm that object coverage is complementary across sources rather than uniform.
Action coverage quantifies the diversity of interaction verbs and reflects the richness of manipulation semantics. We evaluate verb diversity per VQA mode, since different modes are designed to emphasize distinct aspects of embodied behavior. Measuring coverage within functional subsets follows standard practice in lexical diversity analysis and enables mode-aware comparison. The verb diversity is calculated as:
$\mathrm{VerbDiv}(m)=\frac{|\mathcal{V}^{\text{verb}}_{m}|}{N_{m}}\times 1000,$ (2)
where $|\mathcal{V}^{\text{verb}}_{m}|$ denotes the number of unique verb lemmas in mode $m$, and $N_{m}$ denotes the number of QA pairs for that mode. The score is reported as the number of distinct verbs per 1,000 VQA pairs. $VerbDiv$ values are summarized into four descriptive ranges: low (<80<80), medium (8080-$120$), high (120120-$160$), and very high ($\geq 160$).
Measured by VerbDiv, action-centric modes including Reasoning, Mechanics, Temporal, and Summary are predominantly very high across domains. Spatial, Trajectory, and Attribute are mostly medium. This separation is consistent across domains and aligns with the intended role of each mode. This supports that verb coverage is mode-specific and controlled, rather than uniformly distributed across annotations.
@@ 4 Methodology
Figure 4: VLA architecture built on PhysBrain. Given an egocentric observation sequence and a language instruction, PhysBrain encodes multimodal context for action generation. (a) PhysGR00T conditions a flow-matching diffusion action expert on the last-layer hidden states of PhysBrain. (b) PhysPI more tightly couples PhysBrain and the action expert by injecting multiple VLM layers via layer-wise cross-attention.
Using the data annotation pipeline proposed in the previous section, we translate embodied experience from egocentric videos into structured supervision suitable for learning an embodied brain. This process yields E2E-3M, a dataset with roughly 3 million VQA cropus. To preserve general-purpose vision–language capability during SFT, we additionally mix an equal-sized subset sampled from FineVision, a large-scale curated vision–language corpus. We then perform supervised fine-tuning (SFT) on base VLMs (e.g., Qwen2.5-VL-7B) using this mixture, resulting in an egocentric-centered VLM backbone (PhysBrain) with improved first-person understanding, reasoning, and planning capabilities. Quantitative results are reported in Sec. 5 (Tab. 1).
With PhysBrain in hand, we study how these egocentric gains transfer to downstream control under standard VLA instantiations. Our goal in this section is not to propose a new VLA architecture, but to evaluate transferability while minimizing confounding factors from additional heuristics or hand-crafted priors. We follow two widely adopted community paradigms, GR00T-style and Pi-style, and keep the action expert lightweight and consistent across both.
We denote an observation (a short egocentric image sequence) as $o_{t}$, the language instruction as $x$, and the VLM parameters as $\phi$. The VLM produces token-level hidden states
@@
PhysGR00T (A GR00T-Style VLA).
We introduce PhysGR00T, which follow the dual-system design in GR00T N1.5 (Bjorck et al., 2025): the VLM plays the role of System 2 to produce high-level multimodal representations, while a Flow-Matching (FM) action expert (Liu, 2022) serves as System 1 to generate continuous actions. Concretely, PhysGR00T uses the last-layer VLM hidden states $\mathbf{Z}_{t}=\mathbf{H}^{L}_{t}$ as the conditioning signal.
@@
PhysPI (A Pi-Style VLA).
We also instantiate a Pi-style VLA, called PhysPI, in the spirit of $\pi_{0}$ (Black et al., 2024), where the VLM backbone is more tightly coupled with the action expert. Instead of only using the last VLM layerlike PhysGR00T, PhysPI conditions the DiT blocks with multiple VLM layers. Let $M$ be the number of transformer blocks in the action DiT; we take the last $M$ VLM hidden states
$\mathcal{Z}_{t}=\big\{\mathbf{H}^{L-M+1}_{t},\dots,\mathbf{H}^{L}_{t}\big\},$ (7)
and inject them layer-wise into the DiT through cross-attention:
$\mathbf{u}^{(i+1)}=\mathrm{DiTBlock}_{i}\big(Q=\mathbf{u}^{(i)},\,KV=\mathbf{H}^{L-M+i}_{t}\big),\quad i=1,\dots,M,$ (8)
where $\mathbf{u}^{(0)}$ is the embedded (noised) action token sequence. The FM training objective remains identical,
$\mathcal{L}_{\mathrm{FM}}=\mathbb{E}\big[\lVert f_{\theta}(\tilde{\mathbf{a}},\tau;\,\mathcal{Z}_{t},\mathbf{s}_{t})-\mathbf{v}\rVert_{2}^{2}\big].$ (9)
The layer-wise conditioning in PhysPI provides a stronger coupling between intermediate VLM representations and the action expert. This allows us to test whether egocentric improvements distributed across VLM layers can be more effectively utilized for control.
@@ 5 Experiment
This section details the experimental setup, benchmarks, and results. We report results from two primary evaluation tracks: (i) evaluating VLM performance in egocentric settings; and (ii) evaluating performance in a robotic simulation environment following VLA fine-tuning.
@@ 5.1.1 Egocentric Understanding Evaluation
To validate egocentric understanding under a fair and leakage-free setting, we evaluate on EgoThink (Cheng et al., 2024), a widely used benchmark for egocentric reasoning built on Ego4D. Since Ego4D is included in the E2E dataset, our training protocol excludes the Ego4D portion when preparing PhysBrain for EgoThink evaluation. PhysBrain is trained only on the non-Ego4D subsets including EgoDex (Lab) and BuildAI (Factory), and is mixed with an equal-scale sample of general-purpose instruction data to preserve general vision–language capability.
Baselines. We primarily compare our method against two categories of baselines: (i) General VLM, which include closed-source models such as GPT-4 and widely-used open-source models (MiniGPT-4-7B, LLaVA-1.5-7B, LLaMA-3.2-11B, and Qwen2.5-VL-7B); and (ii) Embodied Brain, which include VST-RL-7B (Yang et al., 2025a) and RoboBrain2.0-7B (Team et al., 2025) for comprehensive evaluation.
Evaluation. The comparison methods are evaluated through using the released weight for direct inference. Evaluation conditions are standardized across models. All models use the same prompt template and the generation outputs are scored with a single GPT-4o (Hurst et al., 2024b) judging protocol across all EgoThink subtasks. These controls ensure that performance differences reflect model capability rather than data leakage, prompt variation, or inconsistent scoring.
Table 1 summarizes performance on the six EgoThink dimensions (Activity, Forecast, Localization, Object, Planning, Reasoning). GPT-4 achieves the highest average performance, while our PhysBrain achieves sub-optimal performance and consistently outperforms strong open and competitive baselines. The most pronounced improvement is observed on Planning, where PhysBrain substantially exceeds all baselines and also outperforms GPT-4, indicating a clear advantage in translating egocentric observations into executable plans. Importantly, this improvement is achieved without degrading egocentric perception, under strict Ego4D exclusion during training.
@@
Table 1: Results of evaluating the Egocentric Understanding of VLM models with the EgoThink benchmark. We highlight the best results in bold and the second-best results with underline
Method Activity Forecast Localization Object Planning Reasoning Average General VLM GPT-4 (Achiam et al., 2023) 70.5 61.5 88.5 79 35.5 65.3 67.4 MiniGPT-4-7B (Zhu et al., 2023) 50 15.5 59 48 13 32 36.8 LLaVA-1.5-7B (Liu et al., 2024) 39.5 50 74 62 25.5 51 51.2 LLaMA-3.2-11B (Dubey et al., 2024) 33.5 50 59 64 41 48.7 50.4 Qwen-2.5-VL-7B (Bai et al., 2025c) 56.5 54 71.5 64.7 32 60 57.3 Embodied Brain VST-RL-7B (Yang et al., 2025a) 53 56 70.5 67.7 17 63.7 56.2 RoboBrain2.0-7B (Team et al., 2025) 36 49.5 78 61.3 37 52.7 53.1 PhysBrain (ours) 70 53.5 77 65.3 64.5 58 64.3
5.1.2 Complementary Evaluation on E2E Dataset
To further validate the effectiveness and complementary of the proposed E2E dataset, we evaluate Spatial Aptitude Training (SAT) by performing supervised fine-tuning (SFT) on VST using only E2E data, without introducing any SAT-specific training samples. VST serves as the base model, as it is pre-trained on large-scale, high-quality spatial intelligence datasets and thus provides strong priors for static and object-centric spatial reasoning. This setting allows us to assess whether E2E supervision offers complementary benefits, particularly for egocentric and dynamic spatial reasoning, beyond existing spatial intelligence training.
Prior to fine-tuning, VST attains an overall accuracy of 45.33, with particularly low performance on Egocentric Movement (26.09), indicating limited sensitivity to egocentric motion and viewpoint changes. After fine-tuning on E2E dataset, overall accuracy increases to 59.33, while Egocentric Movement improves markedly to 91.30. Moderate gains are also observed on Action Consequence (54.05 → 64.86) and Perspective (39.39 → 48.48), whereas Object Movement remains comparable (39.13 → 34.78) and Goal Aim is unchanged (58.82). These results indicate that E2E supervision yields targeted improvements in egocentric and dynamic spatial reasoning, complementing the static spatial priors of VST and generalizing without task-specific training data.

### v2

@@ 3.1.3 Quality Assurance and Validation Logic
Open-ended generation easily produces errors that are harmful for training supervision. Common failures include references to non-visible hands and incorrect temporal ordering, et al. We therefore introduce the third stage that a rule checker is designed as a validation gate. Samples that fail validation are rejected and sent back for regeneration with a structured error that indicates the violated constraint.
The checker applies three types of constraints designed to filter visual hallucinations. Evidence grounding restricts the generation scope, requiring that mentioned entities are grounded in the episode-level object metadata. Egocentric consistency enforces the correct hand references and prohibits mentions of unseen limbs or contradictory assignments. Mode-specific temporal logic requires explicit temporal connectors and verifies timeline alignment. The generation–validation loop repeats until all constraints are satisfied. To further guarantee the visual fidelity of the resulting dataset, we conducted a rigorous human audit on a random subset, confirming that our logic-verified filtering effectively retains high-quality, hallucination-free supervision.
3.1.4 Structured Egocentric Corpus
Samples that satisfy all validation constraints are retained and compiled into the egocentric VQA supervision dataset. Each entry records the sampled frames, the selected VQA mode, the generated QA pair, and the validation outcome. This design ensures traceability and reproducibility. The supervision produced by the proposed data engine offers structured and logic-verified supervision that encodes action organization and hand–object interaction, completing the translation of egocentric video data into reliable training signals for egocentric planning and interaction reasoning.
@@ 3.2.1 Data Sources and Domain Coverage
The proposed Egocentric2Embodiment Translation Pipeline is applied to generate the E2E-3M, a large-scale human egocentric video corpora collected across three complementary domains: household, factory, and laboratory environments. These corpora captures substantial variation in environmental context, object composition, and interaction patterns, as shown in Fig.5 and Fig.6.
Specifically, Ego4D represents open-world household activities and provides extensive geographic and contextual diversity. BuildAI captures real industrial workflows, emphasizing procedural regularity and dense hand visibility in factory environments. EgoDex focuses on laboratory settings and offers high-resolution egocentric manipulation sequences with fine-grained interaction cues. These sources differ systematically in spatial layout, object distribution, and task structure. The aggregation yields the E2E-3M dataset with complementary coverage across the space of egocentric embodied experience.
Figure 3: Data Distribution Statistics of E2E-3M dataset.
@@ 3.2.2 Diversity Analysis
To evaluate whether the dataset provides sufficiently rich supervision for generalized embodied manipulation, we analyze diversity along two interpretable axes: object coverage and action (verb) coverage, as quantified in Eq. 1 and Eq. 2.
Object Coverage and Environmental Spectrum. Object coverage reflects the breadth of perceptual contexts. We calculate the normalized noun lemma count per domain as:
$\text{ObjectDiv}(s)=\frac{|\mathcal{V}^{\text{noun}}_{s}|}{T^{\text{noun}}_{s}}\times 1000,$ (1)
where $|\mathcal{V}^{\text{noun}}_{s}|$ is the number of unique noun lemmas and $T^{\text{noun}}_{s}$ is the total noun token count in domain $s$. The distribution shown in Fig.3(b) validates that our aggregated dataset spans the complete spectrum from structured to unstructured environments. Specifically, Household data exhibits high object diversity, capturing the long-tail distribution of objects typical in unstructured, open-world scenarios. Factory data shows lower ObjectDiv scores, which accurately reflect the standardized, repetitive nature of industrial workflows where agents interact with a fixed set of tools and parts. Lab settings fall in between, offering a semi-structured middle ground. This distributional shift is not a limitation but a critical feature: it ensures that VLA models are trained on both the rigid procedural constraints of industry and the chaotic variability of daily life, fostering robust generalization.
Action Coverage and Interaction Semantics. Action coverage quantifies the richness of manipulation semantics. We calculate the verb diversity as:
$\text{VerbDiv}(m)=\frac{|\mathcal{V}^{\text{verb}}_{m}|}{N_{m}}\times 1000,$ (2)
where normalization by the number of QA pairs $N_{m}$ accounts for interaction density. As shown in Fig. 3, action-centric modes (e.g., Reasoning, Mechanics, Temporal) consistently maintain high lexical diversity across all domains. Notably, even in the Factory domain where object diversity is lower due to standardization, the Mechanics and Reasoning scores remain high. This decoupling indicates that while the industrial environment is visually structured, the underlying manipulation tasks involve complex, non-trivial procedural logic. This confirms that our dataset provides dense supervision for manipulation behaviors regardless of the environmental structure.
@@ 4 Methodology
Figure 4: VLA architecture built on PhysBrain. PhysVLA conditions a flow-matching diffusion action expert on the last-layer hidden states of PhysBrain.
Using the data annotation pipeline proposed in the previous section, we translate embodied experience from egocentric videos into structured supervision suitable for learning an embodied brain. To preserve general-purpose vision–language capability during SFT, we additionally mix an equal-sized subset sampled from FineVision, a large-scale curated vision–language corpus. We then perform supervised fine-tuning (SFT) on base VLMs (e.g., Qwen3-VL-4B and Qwen3-VL-8B) using this mixture, resulting in an egocentric-centered VLM backbone (PhysBrain) with improved first-person understanding, reasoning, and planning capabilities. Quantitative results are reported in Table 1.
With PhysBrain in hand, we study how these egocentric gains transfer to downstream control under standard VLA instantiations. Our goal in this section is not to propose a new VLA architecture, but to evaluate transferability while minimizing confounding factors from additional heuristics or hand-crafted priors. We follow the widely adopted community paradigm, GR00T-style, and keep the action expert lightweight and consistent.
We denote an observation (the egocentric image sequence) as $o_{t}$, the language instruction as $x$, and the VLM parameters as $\phi$. The VLM produces token-level hidden states:
@@
PhysVLA.
We introduce PhysVLA, which follows the dual-system design in GR00T N1.5 (Bjorck et al., 2025a): the VLM plays the role of System 2 to produce high-level multimodal representations, while a Flow-Matching (FM) action expert (Liu, 2022) serves as System 1 to generate continuous actions. Concretely, the last-layer VLM hidden states $\mathbf{Z}_{t}=\mathbf{H}^{L}_{t}$ are utilized as the conditioning signal.
@@
Table 1: Comparison on EgoPlan and EgoThink benchmarks. The detailed sub-tasks belong exclusively to the EgoThink benchmark, with the overall average reported in the final column.
Method EgoPlan-B1 EgoPlan-B2 EgoThink Benchmark
Acc. Acc. Act. Fore. Loc. Obj. Asst. Nav. Reas. Avg.
General VLM
GPT-4o (Hurst et al., 2024a) 39.5 41.0 73.0 66.0 89.0 78.6 28.0 12.0 66.0 66.4
MiniGPT-4-7B (Zhu et al., 2023) 28.1 24.5 45.5 36.5 61.5 48.0 30.0 12.0 36.7 41.6
LLaVA-1.5-7B (Liu et al., 2024) 27.8 25.4 35.0 43.5 76.0 65.3 33.0 26.0 53.0 51.6
LLaMA-3.2-11B (Dubey et al., 2024) 24.3 25.1 34.0 49.5 57.5 62.7 42.0 22.0 47.7 48.4
Qwen-3-VL-4B (Yang et al., 2025a) 42.2 34.6 63.5 65.0 82.5 72.6 46.0 35.0 71.0 66.7
Qwen-3-VL-8B (Yang et al., 2025a) 44.3 40.5 68.0 66.5 86.0 72.3 41.0 39.0 61.7 65.9
Embodied Brain
VST-RL-7B (Yang et al., 2025b) 40.8 28.7 55.0 56.5 69.5 67.3 15.0 22.0 62.3 56.2
RoboBrain2.0-7B (Team et al., 2025a) 38.6 23.3 35.0 47.0 77.5 60.7 44.0 38.0 52.3 52.8
RoboBrain2.5-8B (Tan et al., 2026) 45.9 45.2 57.5 56.5 81.0 70.3 40.0 28.0 68.3 62.4
PhysBrain-4B (ours) 43.9 39.3 68.0 64.5 85.5 76.3 66.0 44.0 66.0 69.4
PhysBrain-8B (ours) 47.4 46.9 69.0 69.0 86.5 76.0 65.0 42.0 64.0 69.7
@@ 5 Experiment
This section details the experimental setup, benchmarks, and results. We report results from VLM and VLA evaluations.
@@ 5.1.1 Egocentric Understanding Evaluation
For a comprehensive egocentric evaluation, we conduct experiments on three widely-used benchmarks: EgoPlan-Benchmark1(Chen et al., 2024), EgoPlan-Benchmark2(Qiu et al., 2024), and EgoThink (Cheng et al., 2024). These benchmarks cover diverse real-world settings and complementary protocols: EgoPlan utilizes multiple-choice evaluation, whereas EgoThink scores free-form generations via an LLM judge.
EgoPlan-Benchmark1: EgoPlan-Benchmark1(Chen et al., 2024) evaluates the planning capabilities of Vision Language Models in real-world scenarios from an egocentric perspective, simulating human perception. This benchmark includes realistic tasks involving a diverse range of action plans and intricate visual observations. We validate the VLM models using 3,314 VQA pairs from the EgoPlan-Val set, provides a comprehensive assessment of a model’s ability to understand task progress, track current states, and predict the next feasible action. EgoPlan-Benchmark1 presents significant challenges for current VLMs, highlighting substantial room for improvement in human-level task planning.
EgoPlan-Benchmark2: EgoPlan-Benchmark2(Qiu et al., 2024) extends the scope of its predecessor by incorporating a broader variety of real-world scenarios across four key domains: Work, Daily Life, Hobbies, and Recreation, with 24 distinct scenarios. This expanded benchmark includes 1,321 high-quality multiple-choice questions sourced from 1,113 egocentric videos, each representing a real-world task. The dataset emphasizes task planning, requiring models to predict the next action based on task progress, current visual observations, and task goals. EgoPlan-Benchmark2 provides a rigorous evaluation of VLMs’ planning capabilities and highlights the challenges these models face in real-world decision-making, with detailed analysis and instructions to enhance planning performance.
EgoThink: EgoThink(Cheng et al., 2024) evaluates the first-person perspective capabilities of vision-language models (VLMs) across six core areas: Object, Activity, Localization, Reasoning, Forecasting, and Planning. It contains 700 images from 595 egocentric videos annotated with question-answer pairs. GPT-4 is used as an automatic evaluator to grade open-ended responses. The EgoThink benchmark provides a comprehensive assessment of VLMs’ reasoning and planning abilities in real-world scenarios from a first-person viewpoint, revealing key challenges and offering insights for advancing embodied AI.
@@
Table 2: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment. We highlight the best results in bold and the second-best results with underline.
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average VLA Baselines RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 TraceVLA (Zheng et al., 2025) 12.5 16.6 16.6 65.0 27.7 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.1 0.0 16.6 62.5 27.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (Team et al., 2025b) 64.5 65.5 5.5 93.0 57.1 VLM Baselines Qwen2.5-VL-7B (Bai et al., 2025b) 68.7 35.4 25.0 75.0 51.0 Qwen3-VL-4B (Yang et al., 2025a) 87.5 50.0 29.2 54.2 55.2 Qwen3-VL-8B (Yang et al., 2025a) 68.7 38.5 30.2 87.9 56.3 VST-RL-7B (Yang et al., 2025b) 57.7 41.7 16.7 50.0 41.3 RoboBrain2.0-7B (Team et al., 2025a) 30.8 24.7 2.5 93.3 37.8 RoboBrain2.5-8B (Tan et al., 2026) 75.0 55.5 40.1 100.0 67.6 PhysBrain-4B (ours) 90.3 58.3 34.7 80.6 65.9 PhysBrain-8B (ours) 77.8 62.5 34.8 94.8 67.4
Baselines. We primarily compare our method against two categories of baselines: (i) General VLM, which include closed-source models such as GPT-4o and widely-used open-source models (MiniGPT-4-7B, LLaVA-1.5-7B, LLaMA-3.2-11B, Qwen3-VL-4B, and Qwen3-VL-8B); and (ii) Embodied Brain, which include VST-RL-7B (Yang et al., 2025b), RoboBrain2-7B (Team et al., 2025a) and RoboBrain2.5-8B (Tan et al., 2026).
Evaluation. The comparison methods are evaluated through the released weight for direct inference. Evaluation conditions are standardized across models, and all models use the same prompt template as shown in Table 6. These controls ensure that performance differences reflect model capability rather than prompt variation or inconsistent scoring.
In the Egothink benchmark, the generation outputs are scored with a single GPT-4 (Hurst et al., 2024b) judging protocol across all EgoThink subtasks. Table 1 summarizes performance on the six EgoThink dimensions (Activity, Forecast, Localization, Object, Planning, Reasoning). PhysBrain-8B achieves the highest average performance, while our PhysBrain-4B achieves sub-optimal performance and consistently outperforms strong open and competitive baselines. The most pronounced improvement is observed on Planning, where PhysBrain substantially exceeds all baselines, indicating a clear advantage in translating egocentric observations into executable plans.
In the EgoPlan benchmark, we report accuracy on both EgoPlan-Benchmark1 and EgoPlan-Benchmark2 (Table 1). Trained with our E2E-3M supervision, PhysBrain consistently improves over the Qwen3-VL backbone: PhysBrain-8B reaches 47.4/46.9 on B1/B2, outperforming Qwen3-VL-8B (44.3/40.5) by +3.1/+6.4 points; PhysBrain-4B achieves 43.9/39.3, improving over Qwen3-VL-4B (42.2/34.6) by +1.7/+4.7. These gains support that E2E-3M injects physically grounded interaction priors into the base model, leading to stronger egocentric planning.